In Home and Foreign, there are two factors each of production, land and labor, used to produce only one good. The land supply in each country and the technology of production are exactly the same. The marginal product of labor in each country depends on employment as follows:

Number of Workers Employed

Marginal Product of Last Worker

1

20

2

19

3

18

4

17

5

16

6

15

7

14

8

13

9

12

10

11

11

10

Initially, there are 11 workers employed in Home, but only 3 workers in Foreign. Assume now that Foreign limits immigration so that only one worker can move from Home. Calculate how the movement of the one worker affects the income of five different groups.

What is impact on Foreign landowners is to change costs? The costs in Foreign landowners change from ______ Blank 3 to ______ Blank 4.

What is impact on workers who stay in Home? Workers who stay at home benefit as their wage changes from ______ Blank 5 to ______ Blank 6.

What is impact on Home landowners? Landowners in Home are hurt as their costs change from ______ Blank 7 to ______ Blank 8 for each worker employed.

What is impact on the workers who do move? The workers who do move benefit by seeing their wages change rise from ______ Blank 9 to ______ Blank 10.

What is the total production in two countries? The total production in two countries changes from ______ Blank 11 to ______ Blank 12.
Answer #1

1.What is impact on workers who were originally in Foreign? The wage in Foreign changes from 18 to 16 .

2.What is impact on Foreign landowners is to change costs? The costs in Foreign landowners change from 4.5 to 20.

3.What is impact on workers who stay in Home? Workers who stay at home benefit as their wage changes from 10 to 2.

4.What is impact on Home landowners? Landowners in Home are hurt as their costs change from 60.5 to 30 for each worker employed.

5.What is impact on the workers who do move? The workers who do move benefit by seeing their wages change rise from 10 to 16 .

6.What is the total production in two countries? The total production in two countries changes from 222 to 238 .

(Earlier home production 20 + 19 + 18 = 57 and in foreign 20 + 19 +18+ … + 10 = 165 which equals to 57+165=222

New production achieved by both countries later on 20 + 19 + 18 + 17 + 16 + 15 + 14 = 119 which is done by both foreign and home so 119+119=238 , Hence the change of 16 was the overall increased output achieved)